# AutoRev: Multi-Modal Graph Retrieval for Automated Peer-Review Generation

### v2

@@ D.1. Qualitative Assessment
Examples of the user interface used for the Qualitative Assessment can be seen in Figure 4 and Figure 5.
@@
Figure 4. This interface component corresponds to the Qualitative Assessment task described in Section 5.3.1. It marks the beginning of the review presentation. Due to space constraints, only the summary and a subset of the strengths from one review are shown here.:
@@
Figure 5. This interface component corresponds to the task described in Section 5.3.1. In this part, annotators provide a final rating across the four categories and submit their annotation.:
D.2. Factual Validity Assessment
@@
Examples of the user interface used for Factual Validity Assessment can be seen in Figure 6a and Figure 6b.
@@
Figure 6. The user interface for the Factual Validity Assessment task described in Section 5.3.2. On the left (a), the full research paper is presented to the annotator. On the right (b), the annotator rates each point of the review based on its consistency with the source document.:
Appendix E Example of a Generated Review
We present a review generated by our system for the paper: On the Role of Edge Dependency in Graph Generative Models [^11] https://openreview.net/forum?id=LCQ7YTzgRQ
@@ Appendix E Example of a Generated Review
The paper introduces a novel framework for evaluating generative models of graphs, focusing on the concept of “overlap” to ensure both accuracy and edge diversity. It categorizes graph generative models into three levels of complexity: edge independent, node independent, and fully dependent models. The paper derives theoretical bounds on triangle counts and other short-length cycles based on overlap and provides new generative models for each level. However, the paper’s novelty and practical contributions are questioned due to the use of existing methods and the lack of significant empirical validation. The experimental section is seen as limited and does not convincingly demonstrate the superiority of the proposed methods over existing ones.
@@
  • The paper introduces a novel evaluation framework that considers the importance of model-generated graph overlap to ensure both accuracy and edge diversity, which is a significant contribution to the field of graph generative models.
  • The paper presents a hierarchy of graph generative models categorized into three levels of complexity, which encapsulates a wide range of prevalent methods and provides a structured framework for evaluating graph generative models.
  • The theoretical bounds on the number of triangles and other short-length cycles are well-presented and support the findings with instances demonstrating the asymptotic optimality of these bounds.
  • The authors introduce new generative models for each of the three hierarchical levels, leveraging dense subgraph discovery (Gionis & Tsourakakis, 2015), which is a practical application of the theoretical findings.
@@
  • The novelty of the paper is limited as it primarily relies on existing methods and theoretical frameworks, and the experimental results do not convincingly demonstrate the superiority of the proposed methods over existing ones.
  • The empirical evaluation is weak and does not provide sufficient evidence to support the claims made in the paper. The experimental results are not robust and the paper does not provide enough comparisons with other models to establish the effectiveness of the proposed methods.
  • The paper has several presentation issues, including unclear definitions and notations, which could hinder the understanding of the content for readers not familiar with graph generative models.
  • The practical applications and the significance of the proposed methods are not clearly demonstrated, and the paper lacks a detailed discussion on the limitations and potential future work.
@@
  • Can the authors clarify the definitions and notations used in the paper, particularly the ones related to the overlap and triangle density, to ensure a clearer understanding of the content?
  • How do the authors justify the novelty of their work given the reliance on existing methods and theoretical frameworks?
  • What specific contributions does the paper make to the field of graph generative models, and can these be clearly articulated?
  • Why was the empirical evaluation conducted in a limited manner, and how can the paper be improved to provide more robust and convincing experimental results?
  • Could the authors discuss the practical applications of the proposed methods and the significance of their work in more detail?
  • Are there any plans to extend the theoretical bounds to higher-order cycles, and how might this affect the overall conclusions drawn in the paper?

### v3

@@ F.1 Qualitative Assessment
Examples of the user interface used for the Qualitative Assessment are shown in Figure 5 and  6.
@@
Figure 5: This interface component corresponds to the Qualitative Assessment task described in Section 6.3.1. It marks the beginning of the feedback presentation. Due to space constraints, only the summary and a subset of the strengths from one feedback are shown here.
@@
Figure 6: This interface component corresponds to the task described in Section 6.3.1. In this part, annotators provide a final rating across the four categories and submit their annotation.
F.2 Factual Validity Assessment
Dataset Feedback Positive Rating (%)
ICLR 2024 ICT_10 76.97
Llama $(5,3)$ 65.25
Llama $(5,5)$ 64.58
Other Llama $(5,3)$ 61.86
Llama $(5,5)$ 64.31
Table 12: Overall Positive Rating Percentage (sum of “Completely Agree” and “Mostly Agree”) for test dataset of ICLR 2024 as well as NeurIPS and COLM 2025 (Other).
@@ F.2 Factual Validity Assessment
Examples of the user interface used for Factual Validity Assessment can be seen in Figures 7a and  7b.
@@
Figure 7: The user interface for the Factual Validity Assessment task described in Section 6.3.2. On the left (a), the full research paper is presented to the annotator. On the right (b), the annotator rates each point of the feedback based on its consistency with the source document.
Appendix G Evaluation Guidelines for Factual Validity Assessment
The evaluation guidelines for the factual validity assessment are shown in Figure 8. It is ensured that the annotators selected are eligible to review papers for the particular conference.
**Instructions**:
To assess a structured feedback and check how accurately it reflects the paper’s content.
**Evaluation Process**:
1. Read the entire paper first.
   - If you need to download the paper for some reason, do ensure that you only download the paper using the provided UI.
2. Read the structured feedback and identify every DISCRETE CLAIM.
   - For the “Summary” section of the review, treat the entire content of the Summary
     as ONE discrete claim.
   - For the “Strengths”, “Weaknesses”, and “Questions” sections of the review,
     treat EACH bulleted point as a separate discrete claim.
3. For each claim, choose the best category from the dropdown.
**Claim Categorization**:
You must rate the Validity Agreement of every claim on a scale of 1–4,
where 1 indicates “Completely Disagree” and 4 indicates “Completely Agree”.
**Completely Disagree (Invalid Claim) [1]**:
- The claim is factually wrong, contradicts the paper, or is hallucinated.
**Mostly Disagree (Largely Invalid Claim) [2]**:
- The claim is mostly incorrect, though it may contain minor truth.
**Mostly Agree (Largely Valid Claim) [3]**:
- The claim is mostly correct, with minor gaps or oversights.
**Completely Agree (Fully Valid Claim) [4]**:
- The claim is entirely accurate.
Figure 8: Guidelines used for human evaluation of factual validity of the structured feedback
Appendix H Example of a Generated Structured Feedback
We present a structured feedback generated by our system in Figure 5
On the Role of Edge Dependency in Graph Generative Models
@@ Appendix H Example of a Generated Structured Feedback
The paper introduces a novel framework for evaluating generative models of graphs,
focusing on the concept of ‘‘overlap’’ to ensure both accuracy and edge diversity.
It categorizes graph generative models into three levels of complexity: edge
independent, node independent, and fully dependent models. The paper derives
theoretical bounds on triangle counts and other short-length cycles based on overlap
and provides new generative models for each level. However, the paper’s novelty and
practical contributions are questioned due to the use of existing methods and the
lack of significant empirical validation. The experimental section is seen as
limited and does not convincingly demonstrate the superiority of the proposed
methods over existing ones.
@@
- The paper introduces a novel evaluation framework that considers the importance of
  model-generated graph overlap to ensure both accuracy and edge diversity, which is
  a significant contribution to the field of graph generative models.
- The paper presents a hierarchy of graph generative models categorized into three
  levels of complexity, which encapsulates a wide range of prevalent methods and
  provides a structured framework for evaluating graph generative models.
- The theoretical bounds on the number of triangles and other short-length cycles are
  well-presented and support the findings with instances demonstrating the asymptotic
  optimality of these bounds.
- The authors introduce new generative models for each of the three hierarchical
  levels, leveraging dense subgraph discovery (Gionis & Tsourakakis, 2015), which
  is a practical application of the theoretical findings.
@@
- The novelty of the paper is limited as it primarily relies on existing methods and
  theoretical frameworks, and the experimental results do not convincingly
  demonstrate the superiority of the proposed methods over existing ones.
- The empirical evaluation is weak and does not provide sufficient evidence to
  support the claims made in the paper. The experimental results are not robust and
  the paper does not provide enough comparisons with other models to establish the
  effectiveness of the proposed methods.
- The paper has several presentation issues, including unclear definitions and
  notations, which could hinder the understanding of the content for readers not
  familiar with graph generative models.
- The practical applications and the significance of the proposed methods are not
  clearly demonstrated, and the paper lacks a detailed discussion on the limitations
  and potential future work.
@@
- Can the authors clarify the definitions and notations used in the paper,
  particularly the ones related to the overlap and triangle density, to ensure a
  clearer understanding of the content?
- How do the authors justify the novelty of their work given the reliance on
  existing methods and theoretical frameworks?
- What specific contributions does the paper make to the field of graph generative
  models, and can these be clearly articulated?
- Why was the empirical evaluation conducted in a limited manner, and how can the
  paper be improved to provide more robust and convincing experimental results?
- Could the authors discuss the practical applications of the proposed methods and
  the significance of their work in more detail?
- Are there any plans to extend the theoretical bounds to higher-order cycles, and
  how might this affect the overall conclusions drawn in the paper?
Figure 9: Output of the generated feedback by AutoRev for the paper On the Role of Edge Dependency in Graph Generative Models [^5] https://openreview.net/forum?id=LCQ7YTzgRQ